could you please help!I did these 2 problems but I don't think they are right espically the swcond one I will give you
1. find the total area of the outside surface of a rectangular soild 2.5 m long and 1.7 m wide and 2m high
2.find the total area of the outside surface whose edge measures 4 1/2 ft

Answer by mukhopadhyay(490)

1. Total surface area = Sum of the area of all six sides;
Area of side 1 = 2.5*2 = 5 sq. meter (there are two sides of this area)
Area of side 2 = 2.5*1.7 = 4.25 sq. meter (there are two sides of this area)
Area od side 3 = 1.7*2 = 3.4 sq. meter (here also, there are two similar sides)
Total surface area = 2(5+4.25+3.4) = 2(13.65) = 27.3 sq. meter.
.............
2. It appears to be a cube.
The total surface area is 6[(4.5)^2] = 121.5 sq feet.